What is the rotational kinetic energy of a 60cm diameter bicycle wheel of mass 4.0kg

when the bicycle is travelling at 4.0m/s. Take the radius of gyration k = 50?


1
Expert's answer
2021-05-30T13:30:04-0400

Rotational kinetic energy of the bicycle wheel can be found as follows:


"KE_r=\dfrac{1}{2}I\omega^2,""KE_r=\dfrac{1}{2}mk^2(\dfrac{v}{r})^2,""KE_r=\dfrac{1}{2}\cdot4\ kg\cdot(0.5\ m)^2\cdot(\dfrac{4\ \dfrac{m}{s}}{0.3\ m})^2=89\ J."